find the derivation
9. 6x = 5 x2+1
Answer #1

fare 6105 x2 flalla 960) ba) (a) bagca)- 9(2) 5 (2) [662] 2 bere 912)= 6x-S (8) = x2 + 1 Now f(4) 2 (1271) 5 (61-5) - (61-5- 612 + 102 +6 fal - - 6x 2 + 10x + 6 (12+12